OpenStudy (anonymous):

A bag contains 5 yellow marbles, 7 pink marbles, and 3 purple marbles. Gordon selects a marble without looking. What is the probability that he selects a yellow marble or a purple marble? A. 2/45 B. 1/15 C. 7/15 D. 8/15

OpenStudy (raden):

P(A or B) = P(A) + P(B) let A is the event of selected for a yellow marble and B for a purple marble. the number of yellow marble = 5, so its change is 5of 15(all mable). therefore P(A) = 5/15, while the number of purple marble = 3, therefore P(B) = 3/15 so, P(A or B) = 5/15 + 3/15 = 8/15